3) In a country A, the monetary base is $1,000. People hold one third of their money in the form of
currency and thus two-thirds as bank deposits. Banks hold a third of their deposits in reserve.
(a) What are the reserve-deposit ratio (rr), currency-deposit ratio (cr), money supply (M), and the money
multiplier (m)?
(b) One day, fear about the banking system strikes the population, and people now want to hold half of
their money in the form of currency. If the central bank of country A does nothing, what is the new
money supply?
Answer #1

(a)

rr = Reserves / Deposit = 1/3 (= 0.33)

cr = Currency / Deposit = (1/3) / (2/3) = 1/2 (= 0.5)

m = (1 + cr) / (cr + rr) = (1 + 0.5) / (0.5 + 0.33) = 1.5 / 0.83 = 1.81

M = Monetary base x m = $1,000 x 1.81 = $1,810

(b)

When people hold half their money as currency, they hold half of their money as deposit.

New cr = (1/2) / (1/2) = 1

m = (1 + 1) / (1 + 0.33) = 2 / 1.33 = 1.50

New money supply = $1,000 x 1.50 = $1,500